OpenStudy (queenrae99):

A 70 kg zebra is traveling 10 km/hr east. What is the zebra’s momentum?

OpenStudy (queenrae99):

I multiplied my mass times my velocity and got 700 but i just dont feel like its correct. also it has kg-m/s as what would be at the end of whatever number I get

OpenStudy (queenrae99):

after lots of converting i got 194.44 would that be correct?

OpenStudy (snowsurf):

momentum is equal to \[p = mv\] \[\frac{ 10 km }{ hr }\frac{ 1hr }{ 3600 s }\frac{ 1000m }{ 1km }\] If you make the conversion right your answer should be right.

OpenStudy (whpalmer4):

\[p=mv\]\[70 \text{ kg} * 10 \frac{\text{ km}}{\text{ hr}}*\frac{1000\text{ m}}{1\text{ km}}*\frac{1\text{ hr}}{60\text{ min}}*\frac{1 \text{ min}}{60\text{ s}}\]\[=70 \text{ kg} * 10 \frac{\cancel{\text{ km}}}{\text{ hr}}*\frac{1000\text{ m}}{\cancel{1\text{ km}}}*\frac{1\text{ hr}}{60\text{ min}}*\frac{1 \text{ min}}{60\text{ s}}\]\[=70 \text{ kg} * 10 \frac{\cancel{\text{ km}}}{\cancel{\text{ hr}}}*\frac{1000\text{ m}}{\cancel{1\text{ km}}}*\frac{\cancel{1\text{ hr}}}{60\text{ min}}*\frac{1 \text{ min}}{60\text{ s}}\]\[=70 \text{ kg} * 10 \frac{\cancel{\text{ km}}}{\cancel{\text{ hr}}}*\frac{1000\text{ m}}{\cancel{1\text{ km}}}*\frac{\cancel{1\text{ hr}}}{60\cancel{\text{ min}}}*\frac{1 \cancel{\text{ min}}}{60\text{ s}}\]\[=70\text{ kg}*10*\frac{1000\text{m}}{60*60\text{ s}} = \frac{700000}{3600}\frac{\text{ kg m}}{\text{ s}} \approx 194.4 \frac{\text{ kg m}}{\text{ s}} \] If you set the problem up like this and cancel the units carefully, you will either get the right answer, or a nonsense unit, which is your clue that your answer is incorrect. For example, in this problem it would have been easy to divide by 1000 where you needed to multiply by 1000 in the conversion from km to m. If you write out the units like this, with that mistake you would be left with units of \[\frac{\text{ kg km}^2}{\text{ m s}}\]which doesn't make any sense.

OpenStudy (whpalmer4):

Also, you should specify that that momentum is in the eastward direction, as momentum is a vector quantity.